how to solve x+y=-7 if y=x+1
Answer by stanbon(75887)

how to solve x+y=-7 if y=x+1
Substitute for "y" and solve for "x"::
x + (x+1) = -7
2x = -8
x = -4
----
Then y = -4+1 = -3
----------
Cheers,
Stan H.
-------------